OpenStudy (anonymous):

x+4y=6 2x-3y=1 solve for x and y

OpenStudy (anonymous):

Multiply the first equation by -2 to get the system: -2x-8y=-12, 2x-3y=1. Add the top equation to the bottom equation to get -11y=-11, divide both sides by -11 to get y=1. Substitute y into the top equation to get 2x+4=6, subtract 4 from both sides to get x=2. x=2, y=1

OpenStudy (anonymous):

i dont get something though :/ if 2x+4=6 it would also equal: 2x=6-4 equals: 2x=2 , which means x=1 please explain how it would equal 2

OpenStudy (anonymous):

He is working a different method from what you did earlier so you do like this (addition, subtraction like grade school 2x-3y=1 -2x-8y=-12 becomes 11y=-11. Solve for y. There are a lot of ways to solve simultaneous equations

OpenStudy (anonymous):

yes thank you, but i mean for x i dont know how to solve it

OpenStudy (anonymous):

no, jprahman made a mistake there, I think he meant , when you substitute y in the first equation you get x+4 =6, x= 2

OpenStudy (anonymous):

The 2x+4=6 was a typo, it should have been x + 4 = 6

OpenStudy (anonymous):

makes sense lol :P thank you so much